# On De Giorgi’s lemma for variational interpolants

### v2

@@ 4.4 An alternative route to equality via the chain rule
${\mathcal{E}}({\widetilde{u}_{\uptau_{1}}}^{-})+\uptau_{1}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{1}}({\widetilde{u}_{\uptau_{1}}}^{-}{-}u^{o})\Big{)}+\int_{0}^{\uptau_{j+1}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})\,.$ (4.37)
Conclusion of the proof: Adding (4.36) for $j=1,\ldots,J{-}1$ and (4.37) we obtain that
${\mathcal{E}}({\widetilde{u}_{\sigma_{*}}}^{-})+\sigma_{*}{\mathcal{R}}\Big{(}\frac{1}{\sigma_{*}}({\widetilde{u}_{\sigma_{*}}}^{-}{-}u^{o})\Big{)}+\int_{0}^{\sigma_{*}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})\,.$
In analogy to (4.33) for $\tau_{J}=\sigma_{*}$ we also have $\Phi_{\sigma_{*}}(u^{o};{\widetilde{u}_{\tau_{J}}}^{-})=\phi(u^{o};\sigma_{*})=\Phi_{\sigma_{*}}(u^{o};\widetilde{u}_{\sigma_{*}})$, such that (4.31) is established.
Acknowledgements.
The research of A.M. has been partially supported by Deutsche Forschungsgemeinschaft (DFG) through the Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689) subproject “DistFell”. R.R. acknowledges the support of GNAMPA (INDAM), and of the MIUR-PRIN Grant 2020F3NCPX “Mathematics for industry 4.0 (Math4I4)”.

### v3

@@ 4.5 An alternative route to equality via the chain rule
${\mathcal{E}}({\widetilde{u}_{\uptau_{1}}}^{-})+\uptau_{1}{\mathcal{R}}\Big(\frac{1}{\uptau_{1}}({\widetilde{u}_{\uptau_{1}}}^{-}{-}u^{\circ})\Big)+\int_{0}^{\uptau_{j+1}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ})\,.$ (4.47)
Step 3: Conclusion of the proof. Adding (4.46) for $j=1,\ldots,J{-}1$ and (4.47) we obtain
${\mathcal{E}}({\widetilde{u}_{\sigma_{*}}}^{-})+\sigma_{*}{\mathcal{R}}\Big(\frac{1}{\sigma_{*}}({\widetilde{u}_{\sigma_{*}}}^{-}{-}u^{\circ})\Big)+\int_{0}^{\sigma_{*}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ})\,.$
In analogy to (4.43) for $\tau_{J}=\sigma_{*}$ we also have $\Phi_{\sigma_{*}}({\widetilde{u}_{\tau_{J}}}^{-})=\phi(\sigma_{*})=\Phi_{\sigma_{*}}(\widetilde{u}_{\sigma_{*}})$, such that (4.41) is established.
Acknowledgments.
The authors are grateful to an anonymous reviewer for several improvements and many helpful remarks. The research of A.M. has been partially supported by Deutsche Forschungsgemeinschaft (DFG) through the Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689) subproject “DistFell”. R.R. acknowledges the support of GNAMPA (INDAM), and of the MIUR-PRIN Grant 2020F3NCPX “Mathematics for industry 4.0 (Math4I4)”.